OpenStudy (anonymous):

xy''+y'=-4 y(1)=1 y'(1)=-8 x>0 find y' and y solutions for the IVP

OpenStudy (anonymous):

sorry for my first comment : y'' = (-4 -y') / x y'' = f(x,y') -> f is independent of y so now we can use the sub : y' = p , y'' = dp/dx : dp/dx = (-4-p)/x -dp / (4+p) = dx/x integrate ln(4+p) = -ln(x) + C p = A/x - 4 so y = Aln|x| -4x + B

OpenStudy (anonymous):

now you can find A and B using the conditions.

OpenStudy (anonymous):

Thank you!